OpenStudy (anonymous):

Solve the equation log6(x+5)+log6x=2

OpenStudy (anonymous):

I know that because of the quotient property you would get this log6{x(x+5)}=2 but i'm confused on what the next step would be:/

OpenStudy (anonymous):

log(a)+log(b)= log(A*B)

OpenStudy (anonymous):

I mean product property

OpenStudy (anonymous):

so log (x+5)+log(x)=2 log((x+5)(x))=2 log(x^2+5x)=2

OpenStudy (anonymous):

log(x(x+5))=2

Nnesha (nnesha):

need help ?? :)

OpenStudy (anonymous):

Yes:/ I'm still a little confused

Nnesha (nnesha):

okay after that you have to change log to exponential form for example  right side number became exponent of base and move left side number which is x in my example to the right side